What is the inherent meaning and justification of cross-multiplication as it applies to solving an inequality?
Answer by richwmiller(17219)

Cross multiplication used to be frowned on by teachers because you aren't obviously doing the same thing to both sides.
3/4=x/7
The proper way is to
multiply both sides by 4
3*4/4=x*4/7
3=4x/7
then multiply both sides by 7
3*7=4x*7/7
simplify
3*7=4x
but with cross multiplication we skip the step of showing that we multiply one side by each number.
3/4=x/7
we jump right to the result
3*7=4x
It is shorter and therefore preferred by students.
However there is more chance of errors and not really understanding what you are doing and why.